# Local and Global Trend Bayesian Exponential Smoothing Models

### 03 Feb 2024 v2

@@ 1 Introduction
The ETS models that are currently in use (as described in Hyndman, Koehler, Ord & Snyder (2008), and implemented in the forecast package), are either linear or exponential in their trend, with the only option currently available to potentially bridge this gap being a damped trend. So, the choice is limited especially if the underlying growth rate cannot be suitably modeled by these available options. However, attempts have been made in adjacent fields to develop more flexible trend formulations, for example, Parzen (1982) presents a model where the differencing of ARIMA is replaced by other, slightly more complicated non-stationary AR models, thus allowing for more flexible trend choices. Apart from the relatively unflexible trend, the ETS models assume normality of the errors, homoscedasticity of the errors for additive models, and have some further shortcomings described in details in Section 3.
@@ 1.3 Our Contributions
The methods presented in this paper are available in the Rlgt package on CRAN  [^1] https://cran.r-project.org/package=Rlgt . The Rlgt package is a comprehensive implemention of models discussed in this paper, together with further documentation and illustrations regarding their use.
@@ 3.1 Limited functional form of trend
Despite a variety of different forms of trends being considered in classical ETS models, the choice may frequently still be too limited; this is illustrated by the motivating example of quarterly AWS revenue, shown in Figure 1. Clearly, the trend appears to be growing faster than an additive form but slower than a multiplicative form.
@@ 3.2 Non-normal error distribution
Statistical models based on ETS models, both additive or multiplicative, usually assume that the error terms are identically and independently distributed normal random variables. This assumption is fairly restrictive for most applications, and is not necessarily appropriate for modelling volatile time series data, such as the example from Figure 1. The assumption of normality implicitly assumes that the probability of obtaining a value that highly deviates from the mean (i.e., an outlier) is extremely low. In some real data, such as the movements of stock prices, there is a general consensus that the normality assumption tends to be violated due to high variance, especially in the time of crises. Despite these empirical observations, there are few works in the literature to mitigate this issue. Beaumont (2014) addresses the problem with the use of a Student $t$-distribution instead of a normal distribution. However, such changes are oftentimes not straightforward in the MLE framework and therefore the normality assumption tends to be retained in classical ETS models as even small deviations from the normality assumption may greatly complicate the likelihood function; this can cause considerable issues as maximisation of the likelihood function is the key step in standard MLE-based procedures. Consequently, the assumption of normality is usually required in classical ETS models, as otherwise the MLE may be difficult to find, and other inferential quantities derived from the likelihood, such as confidence intervals may also be very difficult to obtain.
@@ 8.3 Results and comparison
Yearly series
@@
Other series
@@
Monthly series
@@
Quarterly series
@@
Yearly series
@@
Other series
@@
Monthly series
@@
Quarterly series
@@ 8.5 Ablation study and computation times
Yearly series
@@
Other series
@@
Monthly series
@@
Quarterly series
@@
While the LGT and SGT approaches are based on ETS methods, they differ from them in multiple ways. The main differences are concentrated between ETS (AAN) and LGT, while SGT can be considered as a direct extension of LGT into the space of seasonal data; the ETS model closest to SGT is the AAM model. Thus, in this ablation study, for non-seasonal data we investigate differences between ETS (AAN) and LGT to find the importance of the innovations and whether the complete set of these innovations is required to achieve the reported performance. Seasonal data is studied in a similar way. We run performance testing on different variants of our proposed LGT and SGT approaches. The results of the ablation study are reported in Table 7. We remove one by one features such as Student $t$-distribution and heteroscedasticity for errors and replace them with constant variance and normal distribution, respectively. Also, we remove the global trend feature. We note that SGT without global trend has then no trend at all, as SGT has no local trend. As an overall comparison method, we report results of a run of auto ETS (denoted ETS/ZZZ). Furthermore, for non-seasonal data we run our custom Gibbs sampler that is available for the non-seasonal case, and further run ETS/AAN, ETS/AAN (Bayesian version) – ETS (AAN) fitted with Stan using our Bayesian approach, and an implementation of the LGT method within the forecast package in R, which is thus fitted using the standard optimisation method which is used to fit ETS methods. In terms of seasonal data, the forecast package in R does not allow to fit ETS/AAM models due to numerical instabilities, so we fit MAM models instead, using multiplicative error instead of additive error, to compare with the ablation versions of SGT.
@@ 9 Conclusions
The main drawback of our new methods is that they are considerably slower than the original ETS techniques. Though we have been able to overcome this problem to a certain extent by the implementation of a bespoke Gibbs sampler, our methods are still orders of magnitudes slower. However, in the new forecasting world where large scale forecasting is typically performed very competitively by global cross-series models (often Machine Learning models), our method can have an important place in the forecaster’s toolbox to tackle single series forecasting without covariates, in a generally data-scarce environment.

### 22 Mar 2024 v3

@@ 1 Introduction
The ETS models that are currently in use (as described in Hyndman, Koehler, Ord & Snyder (2008), and implemented in the forecast package), are either linear or exponential in their trend, with the only option currently available to potentially bridge this gap being a damped trend. So, the choice is limited especially if the underlying growth rate cannot be suitably modeled by these available options. One way to currently address this shortcoming is by using data transformations as a preprocessing step, such as a logarithm or Box-Cox transformation (Box & Cox, 1964). However, the logarithm is a very strong transformation that can lead to explosive behaviour. The strength of the Box-Cox transformation can be adjusted using its parameter $\lambda$, but choosing this parameter is not trivial and we observe in the experiments of our ablation study that standard methods for this task also often lead to explosive forecasts. Attempts have been made in adjacent fields to develop more flexible trend formulations, for example, Parzen (1982) presents a model where the differencing of ARIMA is replaced by other, slightly more complicated non-stationary AR models, thus allowing for more flexible trend choices. Apart from the relatively unflexible trend, the ETS models assume normality of the errors, homoscedasticity of the errors for additive models, and have some further shortcomings described in details in Section 3.
@@ 1.3 Our Contributions
The methods presented in this paper are available in the Rlgt package on CRAN. [^1] https://cran.r-project.org/package=Rlgt The Rlgt package is a comprehensive implemention of models discussed in this paper, together with further documentation and illustrations regarding their use.
@@ 3.1 Limited functional form of trend
Despite a variety of different forms of trends being considered in classical ETS models, the choice may frequently still be too limited; this is illustrated by the motivating example of quarterly AWS revenue, shown in Figure 1. Clearly, the trend appears to be growing faster than an additive form but slower than a multiplicative form. As mentioned earlier, one way currently available to address this issue is by using a transformation as a pre-processing step, such as a logarithm or a Box-Cox transformation. See, for example, Beaumont (2014). The logarithm is a very strong transform and often leads to explosive behaviour of the forecasts. The Box-Cox transformation has a parameter $\lambda$ that controls its strength, but it is not trivial to choose. Including its estimation in the MLE has the problem that different values of $\lambda$ lead to different autocorrelation structures, making it necessary to, e.g., change the ARIMA order during the search (Guerrero, 1993). The method of Guerrero (1993) chooses $\lambda$ independently of the modelling, in a way that the seasonal variation is stabilised across the series. While in our motivating example of AWS revenue the method is able to achieve a good result similar to our proposed method, in broader experiments, e.g., on the full M3 dataset, the method often still leads to explosive behaviour of the forecasts. See the ablation study in the experimental section of our paper for details.
@@ 3.2 Non-normal error distribution
Statistical models based on ETS models, both additive or multiplicative, usually assume that the error terms are identically and independently distributed normal random variables. This assumption is fairly restrictive for most applications, and is not necessarily appropriate for modelling volatile time series data, such as the example from Figure 1. The assumption of normality implicitly assumes that the probability of obtaining a value that highly deviates from the mean (i.e., an outlier) is extremely low. In some real data, such as the movements of stock prices, there is a general consensus that the normality assumption tends to be violated due to high variance, especially in the time of crises. Despite these empirical observations, there are few works in the literature to mitigate this issue. Beaumont (2014) addresses the problem with the use of a Student $t$-distribution instead of a normal distribution, and a Johnson error trend seasonal transform. However, such changes are oftentimes not straightforward in the MLE framework and therefore the normality assumption tends to be retained in classical ETS models as even small deviations from the normality assumption may greatly complicate the likelihood function; this can cause considerable issues as maximisation of the likelihood function is the key step in standard MLE-based procedures. Consequently, the assumption of normality is usually required in classical ETS models, as otherwise the MLE may be difficult to find, and other inferential quantities derived from the likelihood, such as confidence intervals may also be very difficult to obtain.
@@ 8.3 Results and comparison
Yearly (amount of series: 645, horizon: 6)
@@
Other (amount of series: 174, horizon: 8)
@@
Monthly (amount of series: 1428, horizon: 18)
@@
Quarterly (amount of series: 756, horizon: 8)
@@
Yearly (amount of series: 645, horizon: 6)
@@
Other (amount of series: 174, horizon: 8)
@@
Monthly (amount of series: 1428, horizon: 18)
@@
Quarterly (amount of series: 756, horizon: 8)
@@ 8.5 Ablation study and computation times
Yearly (amount of series: 645, horizon: 6)
@@
Other (amount of series: 174, horizon: 8)
@@
Monthly (amount of series: 1428, horizon: 18)
@@
Quarterly (amount of series: 756, horizon: 8)
@@
While the LGT and SGT approaches are based on ETS methods, they differ from them in multiple ways. The main differences are concentrated between ETS (AAN) and LGT, while SGT can be considered as a direct extension of LGT into the space of seasonal data; the ETS model closest to SGT is the AAM model. Thus, in this ablation study, for non-seasonal data we investigate differences between ETS (AAN) and LGT to find the importance of the innovations and whether the complete set of these innovations is required to achieve the reported performance. Seasonal data is studied in a similar way. We run performance testing on different variants of our proposed LGT and SGT approaches. The results of the ablation study are reported in Table 7. We remove one by one features such as Student $t$-distribution and heteroscedasticity for errors and replace them with constant variance and normal distribution, respectively. Also, we remove the global trend feature. We note that SGT without global trend has then no trend at all, as SGT has no local trend. As overall comparison methods, we report results of a run of auto ETS (denoted ETS/ZZZ), and a run of auto ETS with a Box-Cox transformation as preprocessing. In the Box-Cox transformation, we choose the parameter $\lambda$ with the method from Guerrero (1993), which is also implemented in the forecast package in R. As a deviation from the default behaviour of the implementation, we restrict $\lambda$ to be non-negative, as otherwise we observe an explosive behaviour on some time series and the method returns NA values as forecasts. Furthermore, for non-seasonal data we run our custom Gibbs sampler that is available for the non-seasonal case, and further run ETS/AAN, ETS/AAN (Bayesian version) – ETS (AAN) fitted with Stan using our Bayesian approach, and an implementation of the LGT method within the forecast package in R, which is thus fitted using the standard optimisation method which is used to fit ETS methods. In terms of seasonal data, the forecast package in R does not allow to fit ETS/AAM models due to numerical instabilities, so we fit MAM models instead, using multiplicative error instead of additive error, to compare with the ablation versions of SGT.
@@ 9 Conclusions
The main drawback of our new methods is that the Stan-based methods are orders of magnitude (typically over a thousand times) slower than the original ETS techniques. Though we have been able to overcome this problem to a certain extent by the implementation of a bespoke Gibbs sampler, shaving off one order of magnitude for the non-seasonal case, our methods are still at least hundreds of times slower than the classical ETS methods. However, in the new forecasting world where large scale forecasting is typically performed very competitively by global cross-series models (often Machine Learning models), our method can have an important place in the forecaster’s toolbox to tackle forecasting of a small number of series with low granularity and without covariates, in a generally data-scarce environment.
Acknowledgements
Christoph Bergmeir is supported by a María Zambrano (Senior) Fellowship that is funded by the Spanish Ministry of Universities and Next Generation funds from the European Union. He was also supported by the Australian Research Council under grant DE190100045 for part of the work. Erwin Wibowo undertook the work on the paper while affiliated with Monash University. Alexander Dokumentov was funded by Monash University to undertake this work.